Operations Report – May 07th

Crew 212 Operations Report 07-MAY-2019

SOL: 2

Name of person filing report: Paolo Guardabasso

Non-nominal systems: Pump in loft

Notes on non-nominal systems: The pump is leaking (see previous reports) and today it started turning on automatically every 7 seconds. The Director informed me that this is because the pump tries to compensate the losses due to the leaking.

Generator (hours run last night): 11:15

On (last night): 8:00 PM

Off (this morning): 7:15 AM

Solar— SOC 74% at 6:30 PM

Diesel Reading – 97%

Propane Reading – 46%

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 320 gallons

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 170 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 143310

Toilet tank emptied: No

Deimos rover used: no (out of order)

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 92.6h

Beginning charge:

Ending charge:

Currently charging: Yes

Opportunity rover used: No

Hours: 58.7h

Beginning charge:

Ending charge:

Currently charging: No

Curiosity rover used: No

Hours: 94.1h

Beginning charge:

Ending charge:

Currently charging: Yes

Notes on rovers: Performed an Engineering EVA this morning to unplug Opportunity and plug Spirit in preparation for the afternoon EVA (that got cancelled because of the weather).

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report

CrewCar used and why, where? Off Campus

General notes and comments: an EVA was cancelled because of the weather.

Summary of internet: Download of astronomy images required a high amount of data.

Summary of suits and radios: 4 radios don’t work and suit 1 and 6 have only one functioning fan.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Repair and assembling of drones for tomorrow’s EVA.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report